OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. express your answer in the form y=mx+b (4,9); x+6y=7

OpenStudy (anonymous):

so if it is going to be parallel, then is will have the same slope, just a different y-intercept; First, we must compute with the given points: So 9=6(4)+x? 9 = 24 - 15, right? so the new equation will be y=6x - 15; They're parallel, and it passes through the point (4, 9).

OpenStudy (cwrw238):

the slope of the given line can be found by wtiting it in the slope - intercept form y = mx+c: y = -(1/6)x + 7/6 using the standard form y-y1 = m(x-x1): y-9 = (-1/6)(x - 4) y = (-1/6)x + 2/3 + 9 y = (-1/6)x + 29/3

OpenStudy (cwrw238):

another form of this equation is 6y + x = 58